### v1

@@ 3.3 Uncertainty discussion
The uncertainties in our stellar parameters are small. As a result of the precise analysis that is possible for solar-twins, the mean effective temperature uncertainty of our sample is $10$ K, the mean surface gravity uncertainty is $0.01$, and the mean metallicity uncertainty is $0.01$ dex. Propagating these uncertainties into our spectral synthesis analysis is difficult, as these small changes in the photospheric parameters cause virtually zero difference in the Be fitting. When attempting to quantify it, the uncertainty in beryllium abundance is between 0.01 to 0.02 dex, which is not differentiable in our fitting process.
On the other hand, there are several unknown uncertainties that affect our results. For example, the identification of several transitions in the UV region, and their associated atomic information. This is made clear by our methodology, in which we adjusted line information to match the observed solar spectrum. There are also normalization uncertainties, which can be significant, as no continuum region can be clearly identified. Therefore, a more reliable way to analyze the uncertainty budget in the beryllium abundance determination is directly via the spectral fitting itself. An example is seen in Figure 2, where the fitting of the solar beryllium is shown, along with a typical ±$\pm 0.1$ uncertainty budget. By deriving uncertainties directly via the fitting process, we found the mean abundance uncertainty for the stars in our sample to be ±$\pm 0.12$ dex, which precisely matches the quoted uncertainty by Boesgaard et al. (2022). In their paper, Boesgaard et al. (2022) also argues that the unknown uncertainties in the fitting process of the beryllium abundance surpasses the formal uncertainties due to stellar parameters.
@@
Designation $T_{\rm eff}$a $\log g$a $\rm{[Fe/H]}$a Massa Agea A(Li) A(Be)b
@@ 4 Discussion
The lithium abundances of the stars in our sample have been previously derived in different studies, in particular by Carlos et al. (2019) and Martos et al. (2023). We re-derived all abundances in an effort to homogenize our analysis with the stellar parameters we found, and for the stars in common our abundances reproduce the previously determined ones. For those, we adopt the abundances from the original studies. In Table 2 we show the lithium abundances and their sources. In Figure 3 we show the lithium abundances as a function of age. Our abundances are displayed in purple and the gray squares represent the measurements from Boesgaard et al. (2022) for the stars in their sample that satisfy our solar twin criteria. The dashed line indicate the linear fit to our data (excluding the Boesgaard et al. (2022) stars). Unsurprisingly, as we are using the same linelist and methodology for deriving lithium as previous works, we find good agreement between our abundances and those from previous studies. Although our sample is smaller, and we do have two stars with low lithium abundances (HIP 93858, and HIP 116937), the relationship between lithium abundance and stellar age we find is in agreement with Carlos et al. (2019) and Martos et al. (2023)(A(Li)$=-0.230\times Age+2.39$, or Age$=-3.02\times A(Li)+9.52$, with ages in Gyr). Therefore, our selected sample also display the standard behavior of main-sequence depletion of lithium in solar-twins.
@@
In Figure 4, we show the abundances of beryllium as a function of stellar age. We also include, as gray squares, a sample of stars from Boesgaard et al. (2022) that satisfy our solar-twin criteria. Boesgaard et al. (2022) found a solar abundance of A(Be) $=1.23$ dex in their analysis of the Sun (using a sky spectrum). They performed a 1D LTE analysis using the same radiative transfer code we used (MOOG). There is good agreement between their analysis and the 1D LTE analysis performed by Takeda et al. (2011). We attribute the $0.09$ dex difference between our adopted solar abundances to the inherently different analysis methodologies. While Boesgaard et al. (2022) analyzed their spectra from scratch, we chose to anchor our analysis by the solar abundance found by the most-recent (non-LTE) solar analysis of Korotin & Kučinskas (2022), which is in better agreement with meteoritic beryllium abundance of A(Be) $=1.31$ dex. For more details on the differences and a careful discussion of the NLTE methodology, we refer the reader to Korotin & Kučinskas (2022).
@@
Figure 4: Beryllium abundances of our sample of solar-twins as a function of stellar age. The Sun is plotted as a black solar standard symbol, and the gray squares are the beryllium abundances of stars from Boesgaard et al. (2022) that satisfy our solar-twin criteria, shifted by $0.09$ dex. Our data do is consistent with a flat beryllium distribution as a function of stellar age, consistent with no beryllium depletion.
Therefore, because our linelist, and abundance determination, was designed to reproduce the Korotin & Kučinskas (2022) solar value, we expected there could be a systematic difference between our Be inferences and those of previous works, as our calibration should include NLTE corrections for solar-twins. To account for this systematic difference, we added $0.09$ dex (the difference between our inferred solar abundance to theirs) to all the abundances of the Boesgaard et al. (2022) study that are shown in Figure 4. What we see is that, with this “correction”, our abundances and Boesgaard et al. (2022) abundances are on the same scale.
@@
There is a total of 34 stars that meet our criteria in Boesgaard et al. (2022), and there is an abundance dispersion of $0.11$ dex if we exclude the outliers. In our sample, composed of 23 stars, the abundance dispersion is $0.05$ dex. We attribute the difference in scatter to the higher S/N of our data and to our calibration based on the solar spectrum. Our scatter is comparable to the $0.04$ dex scatter found by Tucci Maia et al. (2015), which used a similar methodology. However, their sample is considerably smaller (8 stars). Tucci Maia et al. (2015) found beryllium abundances that are considerably higher (up to A(Be)$=1.52$ for one of their target stars). We do not include any of their measurements, as there seems to be too large disagreements, which we cannot explain, compared to our data and to other high precision studies of beryllium (e.g., Takeda et al., 2011; Boesgaard et al., 2022).
@@
We fit a linear regression model (dashed black line) to our beryllium abundances as a function of age and find A(Be) $=0.003\times Age+1.28$, or Age $=7.13\times A(Be)-4.42$, for ages in Gyr. Although the slope found in Boesgaard et al. (2022) data is slightly higher than ours ($0.01$ vs $0.003$), the correlation between Be abundance and age is flat, and there is not any discernible beryllium depletion in solar-twins through their main-sequence lifetime. Our results cannot probe pre-main-sequence Be burning, which would require a different observational approach, but we show that the mixing mechanism responsible for dredging lithium and burning it during the main-sequence is unable to dredge material deep enough in the star to reach temperatures capable of Be burning ($3.5\times 10^{6}$ K).
@@
To show how current standard models behave when it comes to beryllium and lithium burning, we did a very simple experiment with the open-source 1D stellar evolution code MESA (Modules for Experiments in Stellar Astrophysics Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023). Following Sevilla et al. (2022), we modified the initial metallicity (Z) in the MESA inlists they used to predict lithium evolution in solar-type stars. Our simple model was calibrated to match present-day solar properties (bolometric luminosity, radius, and temperature) by adopting initial Z=0.016 to account for diffusion effects. Also, to reproduce the secular lithium evolution trend in solar-type stars the model, in its non-rotating version, includes thermohaline mixing (Brown et al., 2013), and elemental diffusion (Paxton et al., 2018) only after ZAMS. For more details on the model, we refer to Sevilla et al. (2022). For beryllium, we used the same toy inlists, but we assumed the protosolar Be abundance of A(Be) $=1.40$ (Lodders, 2021). We show the result of this simple exercise as the solid green line in Figure 4. While the data shows that the initial beryllium abundance at the zero-age-main-sequence (ZAMS) should be the same as the current abundance (A(Be)$=1.32$), our toy model did not include ZAMS burning. Therefore, it only reaches the current value at approximately 7 Gyrs. In our toy model, in the 10 Gyr range probed by our data, beryllium decreases from A(Be)$=1.40$ to A(Be)$=1.23$. Even if we account for our data scatter and uncertainties, the data does not show a $0.17$ dex decrease in abundance. What our simple exercise shows is that, indeed, current models cannot explain what is observed in lithium and beryllium abundances in solar-like stars.
@@
As previously pointed-out, Boesgaard et al. (2022) discussed how their data compares to predictions from models favoring rotational mixing as the mechanism to deplete lithium and beryllium, but the data does not support it. Rotational models deplete too much, up to $0.4$ dex, beryllium during the main sequence (Deliyannis & Pinsonneault, 1990), and the data does not support such level of depletion. Furthermore, these rotational models (e.g. Deliyannis & Pinsonneault, 1997) show a slope between A(Be) and A(Li) of about 0.4 for F stars, with a decrease for cooler stars, such as those in our sample. In Figure 5 we show the relation between beryllium and lithium in our sample. We found a negligible slope of $0.01$. In their study Boesgaard et al. (2022) also found a negligible slope of $-0.008$. Neither slopes are significant and they agree within uncertainties. This is further evidence that rotationally-induced mixing models burn significantly more Be than observed.
@@
We find that our results, for $1\ \mathrm{M_{\odot}}$ stars, are better represented by the convective overshooting models of Xiong & Deng (2007). Their models reproduce the typical lithium burning while no beryllium burning is observed for their $1\ \mathrm{M_{\odot}}$ models during the main sequence (see their Figure 1). The convective settling models by Andrássy & Spruit (2015) also presents itself as a good alternative, as it also does not burn as much beryllium at the temperature range discussed in this paper. As discussed in Boesgaard et al. (2022), mixing models that include gravity waves can also be a good explanation, as they do not burn Be as much as Li.
@@
Interestingly, we do find a clear trend with metallicity even in our very narrow metallicity range ($-0.15\leq\rm{[Fe/H]}\leq+0.15$), where A(Be)$=0.26\times\rm{[Fe/H]+1.29}$, a very discernible slope of $0.26$. We show this result in Figure 6. This slope is comparable to the slope of $0.34$ found by Boesgaard et al. (2022), although their sample extended to metallicities as low as [Fe/H]$=-0.49$ dex.
@@
The only beryllium isotope that is observable in stellar photospheres is 9Be. This isotope cannot be produced in stellar interiors, and its nucleosynthesis is limited to spallation reactions (i.e., breaking apart of heavier nuclei by high-energy cosmic rays). Therefore, as there is no observable depletion of beryllium as a function of age, our results are an excellent tracer of spallation nucleosynthesis for models of the galactic chemical evolution. Our results corroborate the conclusions by Boesgaard et al. (2022).
@@ 5 Conclusion
In this paper we describe the observations and analysis of lithium and beryllium abundances in a sample of solar-twins with a comprehensive age range. We design a new and improved linelist (available upon request) for 1D LTE analysis of beryllium abundances that reproduce the most recent non-LTE Be abundance measurements.
@@
As previously found, we see an increase in the beryllium abundances as a function of metallicity, possibly indicative of galactic beryllium production via cosmic ray spallation.
@@ Acknowledgments
Based on observations collected at the European Southern Observatory under ESO programme 111.24VX. Henrique Reggiani acknowledges the support from NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. Jhon Yana Galarza acknowledges support from a Carnegie Fellowship. Diego Lorenzo-Oliveira acknowledges the support from CNPq (PCI 301612/2024-2). Sofia Covarrubias acknowledges support from the Carnegie Astrophysics Summer Student Internship Program (CASSI) and the Ralph M. Parsons Foundation. J.C. acknowledges support from the Agencia Nacional de Investigación y Desarrollo (ANID) via Proyecto Fondecyt Regular 1231345, and from ANID BASAL project CATA2-FB210003. We also thank Joshua D. Simon and Andrew McWilliam for the (many) useful discussions about this project.

### v2

@@ 3.4 Uncertainty discussion
The uncertainties in our stellar parameters are small. As a result of the precise analysis that is possible for solar-twins, the mean effective temperature uncertainty of our sample is $10$ K, the mean surface gravity uncertainty is $0.01$, and the mean metallicity uncertainty is $0.01$ dex. Propagating these uncertainties into our spectral synthesis analysis is difficult, as these small changes in the photospheric parameters cause virtually zero difference in the Be fitting. When attempting to quantify it, the uncertainty in beryllium abundance is between $0.01$ and $0.02$ dex, a value not discernible by our fitting process. The uncertainties due to our choice of macroturbulent velocity are similarly small. dos Santos et al. (2016) estimated a typical macroturbulent velocity uncertainties of $\sim 0.1$ km.s-1 for solar-twins. Therefore, we do not account for this source of error given its negligible impact on measured Be abundances, given the resolution and SNR of our data.
On the other hand, there are several unknown uncertainties that affect our results. For example, the identification of several transitions in the UV region, and their associated atomic information. This is made clear by our methodology, in which we adjusted line information to match the observed solar spectrum. There are also normalization uncertainties, which can be significant, as no continuum region can be clearly identified. Therefore, a more reliable way to analyze the uncertainty budget in the beryllium abundance determination is directly via the spectral fitting itself. An example is seen in Figure 2, where the fitting of the solar beryllium is shown, along with a typical ±$\pm 0.1$ uncertainty budget. By deriving uncertainties directly via the fitting process, we found the mean abundance uncertainty for the stars in our sample to be ±$\pm 0.12$ dex in excellent agreement with the quoted uncertainty by Boesgaard et al. (2022). In terms of total error budget in Be abundance determinations, the authors claim that unknown sources of uncertainty may surpass the typical impact of stellar parameter errors.
@@
Designation $T_{\rm eff}$a $\log g$a $\rm{[Fe/H]}$a Massa $\tau_{\mathrm{iso}}^{a}$ A(Li) A(Be)b
@@ 4 Discussion
The lithium abundances of the stars in our sample have been previously derived in different studies, in particular by Carlos et al. (2019) and Martos et al. (2023). We re-derived all abundances in an effort to homogenize our analysis with the stellar parameters we found. In Figure 3 we show our lithium abundances as a function of age. Our abundances are displayed in purple and the gray squares represent the measurements from Boesgaard et al. (2022) for the stars in their sample that satisfy our solar twin criteria. The dashed line indicates the linear fit to our data (excluding the Boesgaard et al. (2022) stars). Unsurprisingly, as we are using the same linelist and methodology for deriving lithium as previous works, we find good agreement between our abundances and those from previous studies. The scatter we found in our lithium abundances is quite tipycal of other studies analyzing lithium in solar twins (e.g., Carlos et al., 2016; Rathsam et al., 2023; Martos et al., 2023). The linear correlations between lithium and stellar ages we find are
$\mathrm{A(Li)}=-0.20\pm 0.05\times\tau_{\mathrm{iso}}+2.39\pm 0.26$ (1)
$\tau_{\mathrm{iso}}=-3.02\pm 0.46\times\mathrm{A(Li)}+9.52\pm 0.76$ (2)
with ages ($\tau_{\mathrm{iso}}$) in Gyr. Although our sample is smaller, and we do have two stars with low lithium abundances (HIP 93858, and HIP 116937), Equations 1 and 2 agree with previous works from Carlos et al. (2019), Martos et al. (2023), and Rathsam et al. (2023). The latter found $\mathrm{A(Li)=-0.27\pm 0.02\times\mathrm{\tau_{iso}}+2.65\pm 0.11}$. If we remove the two low lithium outliers (HIP 93858, and HIP 116937), the relationship between lithium and age we find
$\mathrm{A(Li)}=-0.23\pm 0.03\times\tau_{\mathrm{iso}}+2.64\pm 0.19$ (3)
agrees even better with the above mentioned studies. However, we found somewhat higher uncertainties in our fits, due to our reduced sample size. It is important to highlight that lithium in our selected sample has the standard behavior of main-sequence depletion in solar-twins.
We briefly mention our outliers. The star with the highest lithium abundance for its age (HIP 54287 - A(Li)$=1.91\pm 0.01$, with an age of $\tau_{\mathrm{iso}}=6.7\pm 0.9$ Gyr), had already been previously identified as an outlier by Carlos et al. (2019). The second one, HIP 54582 (A(Li)$=1.64\pm 0.02$, with an age of $\tau_{\mathrm{iso}}=6.8\pm 0.4$ Gyr), has a detected brown dwarf companion of 20.7 MJUP (Feng et al., 2022). The $\sim$3.5 Gyr-old stars showing very low lithium content are candidates for a further investigation in the near future since they might resemble the recently reported young solar-twin with the lowest lithium abundance detected so far, HIP 8522 (Yana Galarza et al., 2024).
@@
In Figure 4, we show the abundances of beryllium as a function of stellar age. We also include, as gray squares, a sample of stars from Boesgaard et al. (2022) that satisfy our solar-twin criteria. Boesgaard et al. (2022) found a solar abundance of A(Be) $=1.23$ dex in their analysis of the Sun (using a sky spectrum), very similar to the abundance adopted in this work. There is good agreement between their analysis and the 1D LTE analysis performed by Takeda et al. (2011) as well.
@@
Figure 4: Beryllium abundances of our sample of solar-twins as a function of stellar age. The Sun is plotted as a black solar standard symbol, and the gray squares are the beryllium abundances of stars from Boesgaard et al. (2022) that satisfy our solar-twin criteria. Our data do is consistent with a flat beryllium distribution as a function of stellar age, consistent with no beryllium depletion.
@@
There is a total of 34 stars that meet our criteria in Boesgaard et al. (2022), and there is an abundance dispersion of $0.11$ dex if we exclude the outliers. In our sample, composed of 23 stars, the abundance dispersion is $0.05$ dex. We attribute the difference in scatter to the higher S/N of our data and to our calibration based on the most recent available solar abundances. It is important to say, though, that the differences are small and within measurement uncertainties. Our scatter is comparable to the $0.04$ dex scatter found by Tucci Maia et al. (2015), which also did a differential study. However, their sample is considerably smaller (8 stars). Tucci Maia et al. (2015) found beryllium abundances that are considerably higher (up to A(Be)$=1.52$ for one of their target stars). We do not include any of their measurements, as there seems to be too large disagreements, which we cannot explain, compared to our data and to other high precision studies of beryllium (e.g., Takeda et al., 2011; Boesgaard et al., 2022).
We fit a linear regression model (dashed black line) to our beryllium abundances as a function of age and find
$\mathrm{A(Be)}=0.003\pm 0.004\times\tau_{\mathrm{iso}}+1.18\pm 0.02$ (4)
$\tau_{\mathrm{iso}}=7.13\pm 9.95\times\mathrm{A(Be)}-3.71\pm 11.94,$ (5)
@@
for $\tau_{\mathrm{iso}}$ in Gyr. Although the slope found in Boesgaard et al. (2022) data is slightly higher than ours ($0.01$ vs $0.003$), the correlation between Be abundance and age is flat, and there is not any discernible beryllium depletion in solar-twins through their main-sequence lifetime. Our results cannot probe pre-main-sequence Be burning, which would require a different observational approach, but we show that the mixing mechanism responsible for dredging lithium and burning it during the main-sequence is unable to dredge material deep enough in the star to reach temperatures capable of Be burning ($3.5\times 10^{6}$ K).
@@
Current standard models can’t reproduce beryllium and lithium burning. To check their behavior we did a very simple experiment with the open-source 1D stellar evolution code MESA (Modules for Experiments in Stellar Astrophysics Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023). Following Sevilla et al. (2022), we modified the initial metallicity (Z) in the MESA inlists they used to predict lithium evolution in solar-type stars. Our simple model was calibrated to match present-day solar properties (bolometric luminosity, radius, and temperature) by adopting initial Z=0.016 to account for diffusion effects. Also, to reproduce the secular lithium evolution trend in solar-type stars the model, in its non-rotating version, includes thermohaline mixing (Brown et al., 2013), and elemental diffusion (Paxton et al., 2018) only after ZAMS. For more details on the model, we refer to Sevilla et al. (2022). For beryllium, we used the same toy inlists, but we assumed the protosolar Be abundance of A(Be) $=1.40$ (Lodders, 2021). While the data shows that the initial beryllium abundance at the zero-age-main-sequence (ZAMS) should be the same as the current abundance (A(Be)$\sim 1.21$), our toy model did not include ZAMS burning. Therefore, it only reaches the current value at approximately 10 Gyrs. In our toy model, in the 10 Gyr range probed by our data, beryllium decreases from A(Be)$=1.40$ to A(Be)$=1.23$. Even if we account for our data scatter and uncertainties, the data does not show a $0.17$ dex decrease in abundance over the same period. Therefore, a standard solar evolution model depletes far more beryllium during the main-sequence than what is observed in our sample of solar-twins. What our simple exercise shows is that, indeed, current models cannot simultaneously explain what is observed in lithium and beryllium abundances in solar-like stars.
@@
As previously pointed-out, Boesgaard et al. (2022) discussed how their data compares to predictions from models favoring rotational mixing as the mechanism to deplete lithium and beryllium, but the data does not support it. Rotational models deplete too much, up to $0.4$ dex, beryllium during the main sequence (Deliyannis & Pinsonneault, 1990), and the data does not support such level of depletion. Furthermore, these rotational models (e.g. Deliyannis & Pinsonneault, 1997) show a slope between A(Be) and A(Li) of about 0.4 for F stars, with a decrease for cooler stars, such as those in our sample. In Figure 5 we show the relation between beryllium and lithium in our sample. We found a negligible slope of $0.01$. In their study Boesgaard et al. (2022) also found a negligible slope of $-0.008$. Neither slopes are significant and they agree within uncertainties. This is further evidence that rotationally-induced mixing models burn significantly more Be than observed. Dumont et al. (2021b) and Dumont et al. (2021a) rotationally-induced mixing models for F and G stars at the main-sequence do not find as much Be burning as previous models, but still predict more Be burning than observed in our work (see their Figure 4, and our Figure 5).
@@
We find that our results, for $1\ \mathrm{M_{\odot}}$ stars, are better represented by the models of Xiong & Deng (2007). Their models reproduce the typical lithium burning while no beryllium burning is observed for their $1\ \mathrm{M_{\odot}}$ models during the main sequence (see their Figure 1). The convective settling models by Andrássy & Spruit (2015) also presents itself as a good alternative, as it also does not burn as much beryllium at the temperature range discussed in this paper. As discussed in Boesgaard et al. (2022), mixing models that include gravity waves can also be a good explanation, as they do not burn Be as much as Li.
@@
Interestingly, we do find a clear trend with metallicity even in our very narrow metallicity range ($-0.15\leq\rm{[Fe/H]}\leq+0.15$), where
$\mathrm{A(Be)}=0.26\pm 0.16\times\rm{[Fe/H]+1.19\pm 0.01},$ (6)
a very discernible slope of $0.26$. We show this result in Figure 6. This slope is comparable to the slope of $0.34$ found by Boesgaard et al. (2022), although their sample extended to metallicities as low as [Fe/H]$=-0.49$ dex.
@@
The only stable beryllium isotope that is observable in stellar photospheres is 9Be. This isotope cannot be produced in stellar interiors, and its nucleosynthesis is limited to spallation reactions (i.e., breaking apart of heavier nuclei by high-energy cosmic rays). Therefore, as there is no observable depletion of beryllium as a function of age, our results are an excellent tracer of spallation nucleosynthesis for models of the galactic chemical evolution, corroborating the conclusions by Boesgaard et al. (2022).
@@
Amarsi et al. (2024) present-day solar photospheric beryllium abundance measurement (A(Be)$=1.21\pm 0.05$) is $0.10$ dex lower than the best CI chondrite measurement of A(Be)$=1.31\pm 0.04$ dex (Lodders, 2021). Amarsi et al. (2024) points out that if one assumes the initial solar beryllium abundance to be the one from the meteorites, A(Be)$=1.31$, beryllium has been depleted by $22\pm 11\%$ on top of any effects of atomic diffusion, in tension with Standard Solar Models. Non-Standard Solar Models, on the other hand, predict excessive burning, as we pointed by our work as well. Based on our toy model described above, at the Solar age, the Sun would have depleted $\approx 0.05$ dex of Be, roughly in agreement with the scenario presented in Figure 9 of Amarsi et al. (2024). The Non-Standard model they presented, which includes effects of angular momentum transport, predicts the abundance of A(Be)$=1.21$ at an age of $\sim 1.5$ Gyr. The solar-twins in our sample with estimated ages of $\sim 1.0$ Gyr already have beryllium abundances on the level of A(Be)$=1.21$, which means that any Be burning must have happened at an earlier stage, and quickly, without considerable burning as the star evolves. This Be burning is therefore unrelated to the well-known lithium depletion as a function of time, observed in main sequence solar-twins.
As was also concluded by Amarsi et al. (2024), our work suggests that we still need further modelling developments to study not only Li burning, but Be as well. On the observational side, further work is needed at younger stars, closer to the ZAMS, to constraint at which point Be abundances are depleted to the mean observed value of A(Be)$\sim 1.21$ in the Sun and solar-twins. It is also important to identify the unknown blend at the Be line (Amarsi et al., 2024) and corroborate current best estimate of the solar Be abundance.
@@ 5 Conclusion
In this paper we describe the analysis of lithium and beryllium abundances in a sample of solar-twins with a comprehensive age range. We design a new and improved linelist (available upon request) for 1D LTE analysis of beryllium abundances that reproduce the most recent 3D non-LTE Be abundance measurements.
@@
Our results indicate that the beryllium depletion observed in the sun (from A(Be)$=1.31$ from the meteorites to A(Be)$=1.21$ from the present-day solar photosphere) likely happens before the first Gyr. More work is still needed, at younger stars, to help constrain how this initial burning takes place and why it stops, both from observational and modeling perspectives.
As previously found, we see an increase in the beryllium abundances as a function of metallicity, possibly indicative of galactic beryllium production via cosmic ray spallation and in novaes.
@@ Acknowledgments
Henrique Reggiani acknowledges the support from NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. Jhon Yana Galarza acknowledges support from a Carnegie Fellowship. Diego Lorenzo-Oliveira acknowledges the support from CNPq (PCI 301612/2024-2). Sofia Covarrubias acknowledges support from the Carnegie Astrophysics Summer Student Internship Program (CASSI) and the Ralph M. Parsons Foundation. J.C. acknowledges support from the Agencia Nacional de Investigación y Desarrollo (ANID) via Proyecto Fondecyt Regular 1231345, and from ANID BASAL project CATA2-FB210003. We also thank Joshua D. Simon and Andrew McWilliam for the useful discussions about this project.